OpenStudy (lexiem):

Wren has a 30% chance of being selected as the president of the school sports club and a 80% chance of being elected editor of the school magazine. What is the probability that Wren will be selected as the president of the sports club and will also be elected editor of the school magazine?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (rayaz):

Express the percentages as probabilities Probability that Wren becomes president of the school sports club = \[\frac{30 }{100 }=0.3\] Similarly the probability that he is elected as editor of the school magazine = \[\frac{ 80 }{ 100 } = 0.8\] These two events are independent and the the probability that they both occur = \[0.3\times0.8\] because when two events are independent \[P(A and B) = P(A) \times P(B)\]